OpenStudy (anonymous):

Use a system of equations to solve the following problem: Douglas High School has a total of 89 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many boys play sports at this high school? 54 27 81 62

OpenStudy (jdoe0001):

well, let's see there are 89 boys and girls total b + g = 89 boys are 8 more than twice girls twice girls is => 2g 8 more than THAT is => 2g + 8 so 2g + 8 = b the system of equations looks like b + g = 89 2g + 8 = b so, solve by "substitution" :)

OpenStudy (anonymous):

please help me